### v1

# Intrinsic viscous liquid dynamics

Abstract
When liquids are cooled, their dynamics are slowed, and if crystallization is avoided, they will solidify into an amorphous structure referred to as a glass. Experiments show that chemically distinct glass-forming liquids have universal features of the spectrum and temperature dependence of the main structural relaxation. We introduce Randium, a generic energetically coarse-grained model of viscous liquids, and demonstrate that the intrinsic dynamics of viscous liquids emerges. These results suggest that Randium belongs to a universal class of systems whose dynamics capture the essential physics of viscous liquid relaxation, bridging microscopic molecular models and coarse-grained theoretical descriptions.
preprint: Draft version
I Introduction
Molecular motion becomes slower as liquids are cooled. If crystallization is bypassed, the system solidifies into a disordered structure termed a glass [Dyre2007]. At the glass transition temperature, the viscosity of the liquid becomes so large that it ceases to flow. Various experiments have suggested that chemically distinct glass-forming liquids exhibit generic dynamics in their relaxation spectra and temperature dependence of structural relaxation. This work is directly motivated by recent depolarized dynamic light-scattering investigations by Böhmer, Pabst, Weigl, Helbling, Richter, Gabriel, Zeißler, and Blochowicz [Pabst2021, Bhomer2025], which provide striking evidence confirming a longstanding hypothesis that the spectral shape associated with structural relaxation in molecular liquids can be collapsed onto a common shape [Kohlrausch1854, Cole1941, Davidson1950, Williams1970, Jonscher1977, Jonscher1977, Derrida1980, Derrida1981, Dyre1988, Monthus1996, Bierwirth2017, Zhang2017, Niss2018, Dyre2000, Pabst2021, Denny2003, Dyre2024, Bhomer2025, Hasyim2024, Ong2024, Lam2025, Railton2025]. In this paper, we aim to provide a framework for explaining the generic relaxation of highly viscous liquids.
To this end, we are inspired by results from computer simulations [Schroder2000, Heuer2008, Sciortino2005, Scalliet2022, Berthier2023], which show that, when the glass transition is approached, particles (atoms, molecules, or colloids) are temporarily confined in a cage formed by their neighbors. On a longer time scale, particles may escape the cage in a collective flow event. At low temperatures, these flow events are most likely to reverse. However, in rare events, these fundamental flow events may facilitate other nearby events. A cascade of events will eventually allow the system to flow over a free energy barrier, where the system loses memory of its origin. Due to the separation of time scales, dynamics can be viewed as jumping between local minima in an energy landscape [Goldstein1969, Stillinger1983, Schroder2000]. Thus, the dynamics can be described as a complicated Markov chain of fundamental flow events. We propose the following criteria for a model of the energy landscape of a viscous liquid: i) The thermodynamics of the model should capture the inherent energies of real systems, typically Gaussian [Sciortino2005, Heuer2008, Neophytou2024]; ii) The model should have a sense of space, capturing that fundamental flow events facilitate nearby events; iii) Dynamics should be an intrinsic property, i.e., independent of system size. We conjecture that these characteristics constitute a family of models with universal viscous liquid dynamics. One way to construct such a model is by energetic coarse-graining of an atomistic molecular dynamics simulation. Another is to design a simple model within the family – the approach of this study. Previous studies have already investigated and presented simple models that display some [Sherrington1975, Monthus1996, Derrida1980, Derrida1981, Dyre1988, Hasyim2024, Lam2025, Diezemann1997, Railton2025] or all [Edwards1975, Pedersen2006, Dyre2006, Alba2010, Zhang2017, Vasin2022, Ong2024, Hasyim2024] of these characteristics (discussed later).
The question is: Can a simple model capture the physics of molecular systems? To answer this, we propose an idealized model, Randium, and show that it indeed reproduces the generic dynamics of highly viscous glass-forming liquids. Dynamic facilitation and heterogeneity are emergent phenomena in this model. This is strong evidence that Randium belongs to a family of models that includes the energetically coarse-grained energy landscapes of real molecules. Apparently, this class of systems has similar or identical physics. In the following sections, we will first define the model, present the results, and discuss the findings in the light of other proposed explanations for viscous liquid dynamics.
II Randium
Figure 1: Illustration of the Randium model. The values inside each particle represent the particle type. The color of the line between neighbour particles represents the energy of that type-pair. For clarity, this figure shows an $8\times 8$ lattice while the presented results are for a $192\times 192$ lattice.
Consider a two-dimensional square lattice [Ising1925, Lee1952] with periodic boundary conditions (Fig. 1). Let there be $L$ lattice points in each direction, and populate each point with one particle, so the total number of particles is $N=L^{2}$. Let $(x_{n},y_{n})$ be the position of particle $n$. Assign it the type $m_{n}$ out of a total number of $M$ types giving $\Omega=N!/((N/M)!)^{M}$ microstates if there is a the same number of particles of each type. The energy of a microstate is given as a sum of $2N$ bonds between nearest neighbors in the lattice. In this model, a bond represents a local arrangement of particles. Unlike crystals, the local environments of a typical liquid are manyfold, and from the central limit theorem one expects that local free energies are normally distributed. Specifically, we define an $M\times M$ interaction matrix $I$ where elements are drawn from the standard normal distribution,
$P(I_{uv})=\exp(-I_{uv}^{2}/2)/\sqrt{2\pi},$ (1)
while ensuring that the interaction matrix is symmetric $I_{uv}=I_{vu}$. Here, we use natural units where the standard deviation of the energy distribution is one. The Hamiltonian can then be written as
$H=\sum_{\langle ij\rangle}I_{m_{i}m_{j}}.$ (2)
where $m_{i}$ is the type of the particle at position $(x,y)$ and $m_{j}$ is the type of one of the four nearest neighbors. In the limit where both $N$ and $M$ approach infinity, Randium exhibits trivial Gaussian thermodynamics [Derrida1980, Derrida1981], with an expected energy given by $\langle E\rangle=-2N\beta$ where $\beta$ is the inverse temperature. We note that for real systems the Gaussian is an approximation with a possible cutoff at low energies [Saksaengwijit2004] that may result in an ideal glass state [Kauzmann1948].
Dynamics is defined through Monte Carlo simulations with nearest-neighbor swap attempts, employing Boltzmann’s acceptance criterion. The physical interpretation of a neighbor particle swap is a local fundamental collective motion from one inherent state to another of the fine-grained system. This dynamics ensure that rearrangements are local, and that back jumps are likely. The unit of time is defined as one attempt per particle of the model. Conveniently, the system can be equilibrated at low temperatures with unphysical swaps of particle identities in an ensemble coupled to a bath with a distribution of polydisperse particles, resulting in $N=M$. Both types of dynamics can be implemented using a parallelizable algorithm, allowing for efficient calculations on a graphics processing unit. Below we present results with local particle swaps for a system size of $L=192$ (N=36 864N=36\,864) using between two and 512 independent initial configurations. For this system size, one million swap attempts per particle on an NVIDIA GeForce RTX 4070 take about 5 minutes.
Before continuing our investigation of the properties of Randium, we note that the framework can be generalized to other spatial dimensions, along with corresponding rules for connecting neighboring states. We leave such investigations to future studies.
Figure 2: (a) Overlap order-parameter, $Q(t)$, as a function of time for inverse temperatures ranging from $\beta=0.0$ (dark red) to $\beta=2.0$ (dark blue). A characteristic relaxation time, $\tau$, is defined as where the overlap order-parameter is $\frac{1}{2}$ (gray dashed). At high temperatures (reddish colors), the relaxation is near exponential (black dashed): $\exp(-t/t_{0})$. At low temperatures (blueish colors), the relaxation is closer to a stretched exponential with exponent $\frac{1}{2}$ (red dashed): $A\exp(-\sqrt{t/t_{0}})$. The green-dashed is a high-temperature long-times prediction for $\beta=0$, see Eq. 6. (b) 1Q(t)1-Q(t) on a logarithmic scale.
Figure 3: (a) The overlap order-parameter $Q(t/\tau)$ and (b) $\log(1-Q(t/\tau))$ as a function of scaled time. The orange dashed curve indicates a universal curve that $Q(t)$ approaches at intermediate and long times.
Figure 4: Comparing the relaxation of Randium (orange dashed) with molecules measured by depolarized dynamic light scattering. The agreement is excellent.
Figure 5: Temperature dependence of the relaxation time, $\tau(T)$. The red and green dashed lines is a prediction for the high temperature limit, see Appendix A. The blue dashed line is a parabolic scaling [Elmatad2009], $\tau_{q}\exp(J^{2}[\beta-\beta_{q}]^{2})$, with $\tau_{q}=50(1)$, $J=4.3(1)$, $\beta_{q}=0.93(3)$. By extrapolating, the inverse glass-transition temperature is estimated to $\beta_{g}=2.16$ (defined as $\tau(\beta_{g})=10^{14}$). The inset shows decoupling of two timescales, here half-life $\tau$ and self-diffusion $D$, at low temperatures ($\beta>1$).
Figure 6: Spatial distribution of relaxed regions at $t=\tau$ for a range of $\beta$ values. Black corresponds to a site where the particle type has changed, and white to a site where it is unchanged.
Figure 7: (a) Reduced length scale ($\xi/\xi_{0}$) of dynamical heterogeneity at $t=\tau$ (Fig. 6) estimates as a exponential fit, $\exp(-r/\xi)$, to the ”spin-spin” correlation function, $G(r)=\langle\sigma_{i,j}\sigma_{i,j+r}\rangle$ where $\sigma_{i,j}=-1$ if the site is unchanged, and $\sigma_{i,j}=+1$ otherwise (inspired by analysis of the 2D Ising model [Shrock2022]). $\xi/\xi_{0}$ is shown agains the reduced inverse temperatur, $\beta/\beta_{g}$. (b) The reduced length scale ($\xi/\xi_{0}$) as a function of relaxation time $\log(\tau)$. The black dashed line is a $\tau\propto\exp(\xi)$ fit, suggesting that that $\xi/\xi_{0}=6$ at the glass-transition temperture ($\tau_{g}=10^{14}$).
III Results
To monitor dynamics, we define the overlap, $O(t)$, as the fraction of sites that are occupied by the same particle after a time interval $t$. Let $Q(t)=\langle O(t)\rangle$ be the overlap function averaged over initial configurations. Fig. 2(a) shows $Q(t)$ for inverse temperatures ($\beta$’s) ranging from zero (dark red) to 2.0 (dark blue). At high temperatures, the relaxation is nearly exponential (black dashed), whereas at low temperatures, it follows a stretched exponential with an exponent of $\tfrac{1}{2}$ (red dashed): $A\exp(-\sqrt{t/t_{0}})$, where $A\simeq 0.98$. In the frequency domain, this corresponds to a minimum slope of the main relaxation peak of $\tfrac{1}{2}$, consistent with dielectric experiments [Nielsen2009]. At the lowest temperatures, a plateau develops from particle backjumps, as shown in Fig. 2(b) by plotting $\log(Q(t)-1)$. This explains why $A$ is slightly less than one in an empirical fit to a stretched exponential.
In agreement with experimental results [Niss2018], the dynamics of low temperatures show time-temperature superposition. To show this, we define a characteristic relaxation, $\tau$, as the half-life time, defined as the time where half of the lattice sites (on average) have changed,
$Q(\tau)=\frac{1}{2}.$ (3)
Fig. 3(a) shows that for the lowest investigated temperatures, $Q(t/\tau)$ collapses to a universal relaxation curve (orange dashed). Figure 3(b) show $1-Q(t/\tau)$ on a logarithmic scale. Interestingly, the scale-invariant relaxation curve is not a stretched exponential (compare to the red dashed curve).
How does the shape of the relaxation curve of Randium compare to the generic relaxation of experimental data on molecules? To answer this, we reanalyze depolarized dynamic light scattering data presented in Ref. [Pabst2021, Bhomer2025]. Figs. 4(a) and 4(b) show that the empirical data follow the universal curve of Randium. Elmatad, Chandler, and Garrahan [Elmatad2009] has shown that, at low temperatures, the relaxation-time of molecular systems follows a parabolic scaling, $\tau(T)=\tau_{0}\exp(J^{2}(\beta-\beta_{0})^{2})$ in agreement with Randium, see blue dashed line on Fig. 5. The dynamical range from high-temperature dynamics to the glass-transition for molecular liquids typically spans $15$ orders of magnitude ($10^{-13}$ s at high temperature, to $10^{2}$ s at the glass-transition temperature). From this we estimate the inverse glass-transition temperature of Randium to $\beta_{g}=2.16$ (see ++ on Fig. 5). The Angell fragility index at the glass-transition temperature, here defined as $m\equiv\left.\frac{d\log_{10}\tau}{d[\beta/\beta_{g}]}\right|_{\beta_{g}}$ giving $m=2J^{2}\beta_{g}^{2}(1-\beta_{q}/\beta_{g})/\ln 10$, is 43 – within the range of typical molecular glass-formers (this value will likely be different for lattices with other connectivity).
In summary, Randium reproduces time-temperature superposition, the universal relaxation spectrum and the universal shape of the structural relaxation time ($\tau(T)$) of molecular systems. We refer to as the intrinsic viscous liquid dynamics.
IV Discussion
Why does a simple model, here exemplified with Randium, reproduce the intrinsic viscous liquid dynamics of molecular systems? To answer this, recall that dynamical heterogeneity [SchmidtRohr1991, Ediger2000, Tanaka2025] plays a crucial role in understanding viscous liquid dynamics. Specifically, at low temperatures, there are regions of space where particles relax quickly, and regions where structural changes are more sluggish. This results in dynamical heterogeneity-induced decoupling of timescale (exemplified by Stokes-Einstein breakdown [Fujara1992]), at low temperatures, as reproduced by Randium: The inset in Fig. 5 illustrates this. The panels in Fig. 6 show sites where the particle type changes (black) after time $t=\tau$. Interestingly, as temperature is lowered (increase of $\beta$), the cooperatively rearranging regions increase in size, as suggested by Adam and Gibbs [Adam1965]. Figure 7(a) show that the characteristic length-scale, $\xi$, increase more than a factor of 4 in the investigated temperature range. To a good approximation, the relaxation times scales as $\tau\propto\exp(\xi)$. An extrapolation suggest that at the glass-transition temperature, the length-scale is increased by a factor of 6.
What is the origin of dynamic heterogeneity in Randium? To answer this, imagine a low-temperature configuration where particle neighbours have favorable energies. When two particles swap, each of them will gain three new neighbours that likely have unfavorable energies. Thus, it is most likely that particles will swap back. However, in rare events, the initial swap may facilitate nearby swaps, allowing particles to find four new neighbours with a favorable pair energy. This will involve the rearrangement of a region of particles, creating an area of mobility. This area of mobility may facilitate dynamics in nearby areas since particles in that area now have new possibilities of meeting new neighbours with possible favorable neighbours. This is likely similar to what happens in a molecular liquid, where local rearrangements of molecules can trigger cascades of cooperative motion, leading to regions of high mobility embedded in an otherwise rigid structure.
How does Randium compare to other proposed explanations of generic viscous liquid dynamics? Historically, the first descriptions are empirical approaches such as fits to a stretched exponential [Kohlrausch1854, Williams1970] in the time-domain, or the Cole-Cole fit in the frequency-domain [Cole1941, Davidson1950].
More theoretically founded approaches include kinetic facilitation models [Lam2025], spin-glass models [Edwards1975, Sherrington1975, Derrida1980, Derrida1981, Hopfield1982, Parisi1979, Parisi1980, Parisi1983, Sherrington2025, Dahlberg2025], energetic barrier and trap models [Dyre1988, Bouchaud1992, Monthus1996, Diezemann1997, Monthus2003, Schrder2008], elastic models [Dyre2006, Vasin2022, Hasyim2024], and the recently proposed Hyper-sphere model [Railton2025]. Like many of these approaches, Randium builds on the idea of an intrinsic energy landscape put forward by Goldstein in 1969 [Goldstein1969]. In particular, the distinguishable-particle lattice (DPL) model by Lam and coworkers [Zhang2017, Ong2024, Lee2025], and the lattice-gas on a random energy landscape in three dimensions [Pedersen2006] are closely related. Like Randium, these models are defined as particles on a lattice – unlike Randium, dynamics are defined as particles moving into a void, like in the kids’ toy 15-Puzzle [Ratner1990]. The motivation for this dynamics is string-like motions seen in computational studies of atomic glass-formers [Donati1998, Schroder2000]. In Randium, the swap of two particles is interpreted as a fundamental transition between local minima, which, for the atomic model, are string-like motions of tens of particles [Vogel2004]. Thus, Randium can be viewed as more coarse-grained than the models of Refs. [Pedersen2006, Zhang2017, Ong2024, Lee2025] allowing for a large dynamical range. Dynamical mobile regions of the DPL model are dictated by the locations of voids, in contrast, dynamical heterogeneity is an emergent phenomenon.
V Concluding remarks
We emphasize that Randium is not another toy model, but is grounded in physical insights from atomistic simulations and experiments. Randium is motivated by insights into the inherent energy landscape of off-lattice models of molecular systems. We have shown that Randium successfully reproduces the intrinsic viscous liquid dynamics observed in such systems. This provides strong evidence that Randium belongs to a broader class of models governed by similar physics. We conjecture that this class includes variations of Randium with different connectivity, such as a simple cubic lattice in three dimensions, as well as alternative distributions of bond energies. Importantly, it also encompasses the inherent energy landscapes of molecular systems themselves. The precise realization of a Randium-like system appears to have little influence on the universal dynamical behavior, aside from trivial scaling factors. Finally, since Randium is significantly simpler than the inherent energy landscape of molecular systems, it offers the possibility of connecting to more fundamental, analytically tractable models [Parisi1979, Bouchaud1992, Schrder2008]. In this sense, Randium may serve as a stepping-stone framework, bridging realistic molecular models with highly idealized approaches such as the random barrier model, the trap model, or kinetically constrained models.
Acknowledgements.
The authors thank Thomas Blochowicz, Mark Ediger, Tina Hecksher, Andreas Heuer, Camille Scalliet, Walter Kob, Thomas B. Schrøder, and Jeppe C. Dyre for helpful discussions, and in particular Florian Pabst for curating and discussing experimental data.
Appendix A High temperature dynamics
To make a theoretical prediction for the half-life $\tau$ at infinite temperature ($\beta=0$), we may consider the dynamics of Randium as a lattice gas of non-interacting particles. In each step, two sites out of $N\gg 1$ are selected at random and swapped. Thus, the probability that a given site participates in an update is $\frac{2}{N}$, while the probability that it remains untouched is $1-\frac{2}{N}$. After $k$ steps, the probability that a given site has not yet been updated is
$Q(k)=\left(1-\frac{2}{N}\right)^{k}.$ (4)
For $N\to\infty$ this simplifies to
$Q(t)=\exp(-2t)\quad(t\ll 1)$ (5)
where $t\equiv k/N$ is the definition of time. At long-times a given particle makes a random walk on a square lattice, and $Q(t)$ is given by the return probability of a two-dimensional Gauss-distribution:
$Q(t)=\frac{1}{2\pi t}\quad(t\gg 1)$ (6)
Let $\tau_{0}$ be the time required for half of the sites to remain unvisited (at $\beta=0$), i.e. $Q(\tau_{0})=\frac{1}{2}$. From Eq. (5) we get
$\tau_{0}=\ln(\sqrt{2})\simeq 0.35\quad(\beta=0).$ (7)
To provide a description for the $\beta$ dependence we assume an Arrhenius dependence, $\tau=\tau_{0}\exp(\beta A)$. We find empirically that $A=3/2$:
$\tau=\ln(\sqrt{2})\exp\left(3\beta/2\right)\quad(\beta\to 0),$ (8)
see red dashed line on Fig. 5. A more accurate empirical description is
$\tau=\ln(\sqrt{2})\exp\left(3(\beta+\beta^{2})/2\right)\quad(\beta\to 0).$ (9)
See green dashed line on Fig. 5.